# Benchmarking Large Vision-Language Models on Fine-Grained Image Tasks: A Comprehensive Evaluation

### v2

@@ 5 Concluding Remarks
Limitations.
Despite FG-BMK covers most fine-grained image tasks, it is still impossible to cover every fine-grained scenario in real-world applications, such as medical imaging or remote sensing. Nevertheless, we believe that FG-BMK provides the most practical setting for evaluating LVLMs in most use cases, and the conclusions drawn from this benchmark should be applicable to real-world fine-grained visual tasks.
@@ A.1 Evaluation Task Details
In the cross meta-class classification task, we follow the DINOv2 [33] method to train the model on a unified training set where fine-grained categories from different datasets are combined. The model is then tested on each individual dataset to evaluate its performance.
@@ A.2 Data Curation
Wine [39] Industrial 11 4,516
DeepFashion [26] Clothes 46 18,000
SkinCon [9] Dermatology 48 3,866
Flowers102 [32] Flower 102 7,169
Food101 [7] Food 101 101,000
FGVC Aircraft [29] Aircraft 100 6,667
Stanford Dogs [18] Dog 120 20,580
Stanford Cars [19] Car 196 16,185
CUB-200-2011 [41] Bird 200 11,788
VegFru [14] Vegetable 292 146,131
Products-10K [4] Retail 9,691 197,307
iNat2021 [40] Plants 10,000 2,786,843
@@ Appendix B Evaluated Models
As shown in Table 7, we select nine widely-used open-source LVLMs, two closed-source models (GPT-4o [1] and Gemini-2.0-flash [13]) and one purely visual model, each of which employs a distinctive training recipes, including variations in vision encoder, language model, training losses and data.
Table 7: Configurations of the evaluated models. “DINOv2-L” is a purely visual model. “Con” stands for the contrastive loss, “Gen” for the generative loss, “Mat” for the image-text matching loss, “Rec” for the reconstruction loss as used in BEiT3 [42], and “Dis” for the distillation loss as applied in DINOv2 [33].
@@
  • EVA-CLIP [38] aligns visual and textual features using contrastive loss, leveraging over 2 billion web image-text pairs and advanced optimization techniques.
  • InternVL3 [55] adopts a unified pre-training approach over both multimodal and pure-text data, enhanced by variable visual position encoding (V2PE) and advanced post-training strategies for improved scalability and effectiveness.
  • InternVL [8] leverages contrastive, matching, and generative losses in a multi-stage training process, with a large-scale vision encoder and over 6 billion image-text pairs to align visual and textual representation.
  • BLIP-2 [21] bridges the modality gap between frozen image encoders and LLMs using a lightweight Q-Former, leveraging contrastive, matching, and generative loss in a two-stage pre-training process over 129 million data with fewer trainable parameters.
  • Qwen2.5-VL [3] combines dynamic-resolution Vision Transformer with Window Attention to reduce computational cost while preserving native image resolution.
  • Qwen-VL [2] employs a three-stage training process with generative loss, using a VL adapter to align visual and textual features while reducing computational cost over 1.4 billion image-text pairs.
  • CoCa [50] adopts task-specific attentional pooling to tailor visual representations for different training objectives, applying contrastive loss to train the first half of the decoder and generative loss to train the full decoder in an end-to-end manner over 5 billion image-text pairs.
  • BEIT3 [42] treats images as a foreign language, leveraging a mask-then-predict objective over 36 million image-text pairs to unify vision and language pretraining, and introduces a multiway transformer architecture for general-purpose modeling.
  • LLaVA [24] aligns visual and textual features using a simple MLP with generative loss, leveraging 1.2 million GPT-4  generated multimodal instruction-following data for training.
  • DINOv2 [33] uses a self-supervised learning approach, leveraging knowledge distillation and a mask-then-predict strategy over 142 million images to train the vision encoder.
@@ C.1 Results of Hierarchical Granularity Recognition
Figure 2 shows InternVL3’s [55] accuracy in answering true/false and multiple-choice questions within hierarchical granularity recognition task on CUB-200-2011 dataset. In Figure 11, we present additional results for GPT-4o [1], Gemini-2.0-flash [13], Qwen2.5-VL [3], LLaVA [24] and InternVL [8] on CUB-200-2011 [41] and iNat2021 [40] datasets. As shown in the experiments, the accuracy of all models decreases as the granularity becomes finer. When the granularity level reaches the finest level, the models struggle to distinguish between closely related species.
@@
Figure 11: Results of GPT-4o [1], Gemini-2.0-flash [13], Qwen2.5-VL [3], LLaVA [24] and InternVL [8] on true/false and multiple-choice questions across different levels of granularity on CUB-200-2011 [41] and iNat2021 [40] dataset. The x-axis denotes the granularity of the recognition questions.
@@ C.2 Results of Knowledge Bias Estimation
In Figure 3, we observe that LLaVA exhibit highly inconsistent recognition abilities across categories. We also conduct experiments with Qwen2.5-VL, GPT-4o, and Gemini-2.0-flash on fine-grained datasets such as Aircraft [29], Flowers102 [32] and Stanford Dogs [18]. As shown in Figure 12 and Figure 13, all LVLMs display similar trends, indicating inconsistent recognition abilities across fine-grained categories. However, after fine-tuned on datasets with balanced occurrences of fine-grained categories, LVLMs demonstrate remarkable recognition abilities across all fine-grained categories.
@@
Figure 13: Comparison of the original and fine-tuned Qwen2.5-VL [3] models on occurrence-balanced fine-grained aircraft categories. True/false question accuracy for each category is ranked, with blue dots representing the original model and yellow dots the fine-tuned model.
@@ C.3 Results of Attribute Recognition
Table 8: Attribute recognition accuracy of LLaVA [24] on the CUB-200-2011 [41] dataset (values in parentheses represent the average accuracy for each attribute).
@@
Table 9: Attribute recognition accuracy of BLIP2 [21] on the CUB-200-2011 [41] dataset (values in parentheses represent the average accuracy for each attribute).
@@
Table 10: Attribute recognition accuracy of InternVL [8] on the CUB-200-2011 [41] dataset (values in parentheses represent the average accuracy for each attribute).
@@
Table 11: Attribute recognition accuracy of Gemini-2.0-flash [13] on the CUB-200-2011 [41] dataset (values in parentheses represent the average accuracy for each attribute).
@@
Table 12: Attribute recognition accuracy of Qwen2.5-VL [3] on the CUB-200-2011 [41] dataset (values in parentheses represent the average accuracy for each attribute).
@@ D.1 Qualitative Analysis of Granularity Inconsistency in LVLM Alignment Data
D.2 Results of Classification Across Multi-categories
@@
Figure 15: Classification results of LVLM visual features on fine-grained datasets. “Single" denotes accuracy from training on a single meta-category, while “Multiple" reflects accuracy from training on a unified dataset combining multiple meta-categories.

### v3

@@ 5 Concluding Remarks
Ethics Statement
All data used in this study are collected from publicly available datasets that comply with standard research ethics. The benchmark tasks and annotations in FG-BMK are constructed through automatic or rule-based question generation without involving human subjects. Although the dataset SkinCon relate to medical images, it is fully anonymized and released under research-permissive licenses. Our use strictly follows its terms and involves no patient-identifying information, clinical decision-making, or sensitive personal data. We do not release any private or sensitive information, nor do our methods introduce additional safety, fairness, or legal concerns. Overall, FG-BMK provides a practical and ethically compliant testbed for evaluating LVLMs in fine-grained real-world scenarios.
Reproducibility Statement
We ensure reproducibility by using only publicly available datasets and by clearly describing our benchmark construction process in Section 3 and Appendix A. The evaluation pipeline, including prompt design, scoring methods, and model configuration, is detailed in Section 3, Appendix A and Appendix B. We provide extensive implementation details and dataset statistics in the Appendix. All code, data splits, and evaluation scripts has be released.
@@ A.1 Evaluation Task Details
In the cross meta-class classification task, we follow the DINOv2 (Oquab et al., 2023) method to train the model on a unified training set where fine-grained categories from different datasets are combined. The model is then tested on each individual dataset to evaluate its performance.
@@ A.2 Data Curation
Wine (Tianchi, 2021) Industrial 11 4,516
DeepFashion (Liu et al., 2016) Clothes 46 18,000
SkinCon (Daneshjou et al., 2022) Dermatology 48 3,866
Flowers102 (Nilsback and Zisserman, 2008) Flower 102 7,169
Food101 (Bossard et al., 2014) Food 101 101,000
FGVC Aircraft (Maji et al., 2013) Aircraft 100 6,667
Stanford Dogs (Khosla et al., 2011) Dog 120 20,580
Stanford Cars (Krause et al., 2013) Car 196 16,185
CUB-200-2011 (Wah et al., 2011) Bird 200 11,788
VegFru (Hou et al., 2017) Vegetable 292 146,131
Products-10K (Bai et al., 2020) Retail 9,691 197,307
iNat2021 (Van Horn et al., 2021) Plants 10,000 2,786,843
@@
We also expanded the original template set to 10 diverse human-written prompts and reconstructed the multiple-choice questions in the human-oriented benchmark to examine the potential impact of linguistic diversity. As shown in Table 7 and Table 8, increasing the number of templates leads to only minor changes in accuracy, and the overall LVLM behavior and observed trends remain consistent. Therefore, as long as the template clearly states the question, the effect of the template quantity on the results is negligible.
Table 7: Attribute recognition accuracy of InternVL3 (Zhu et al., 2025) using original and extended prompts on the CUB-200-2011 (Wah et al., 2011) dataset (values in parentheses represent the average accuracy for each attribute). Accuracy are shown in the format “original / extended”, with the left representing accuracy using the original prompt and the right using the extended prompt.
Color Attribute (47.40 / 47.45)
belly color 58.49 / 60.04 back color 34.98 / 36.33 bill color 51.31 / 49.64
breast color 54.25 / 55.91 crown color 55.30 / 54.01 eye color 84.59 / 82.96
forehead color 53.32 / 51.90 leg color 44.01 / 45.67 nape color 39.24 / 38.02
throat color 52.77 / 54.53 under tail color 34.69 / 35.80 underparts color 56.20 / 55.08
upper tail color 37.30 / 38.77 upperparts color 28.75 / 27.50 wing color 30.16 / 31.88
primary color 43.05 / 41.29
Pattern Attribute (50.13 / 50.28)
back pattern 40.94 / 39.38 belly pattern 68.13 / 67.00 breast pattern 65.12 / 66.87
head pattern 35.92 / 34.66 tail pattern 41.64 / 42.93 wing pattern 49.04 / 50.84
Shape Attribute (30.95 / 31.01)
bill shape 37.61 / 36.41 shape 52.37 / 50.60 tail shape 10.42 / 12.04
wing shape 23.39 / 24.98
Length Attribute (71.03 / 69.71) Size Attribute (52.55 / 54.21)
bill length 71.03 / 69.71 size 52.55 / 54.21
Table 8: Results of InternVL3 using original and extended prompts on true/false (TF) and multiple-choice (MC) questions across different levels of granularity on the CUB-200-2011 dataset. Results are shown in the format “original / extended”.
TF
Class 98.79 / 98.03 Genus 85.69 / 86.19 Species 61.88 / 62.34
MC
Class 99.42 / 99.58 Genus 88.13 / 87.62 Species 60.15 / 59.23
@@ Appendix B Evaluated Models
As shown in Table 9, we select nine widely-used open-source LVLMs, two closed-source models (GPT-4o (OpenAI, 2023) and Gemini-2.0-flash (Gemini Team, 2024)) and one purely visual model, each of which employs a distinctive training recipes, including variations in vision encoder, language model, training losses and data.
Table 9: Configurations of the evaluated models. “DINOv2-L” is a purely visual model. “Con” stands for the contrastive loss, “Gen” for the generative loss, “Mat” for the image-text matching loss, “Rec” for the reconstruction loss as used in BEiT3 (Wang et al., 2023), and “Dis” for the distillation loss as applied in DINOv2 (Oquab et al., 2023).
@@
  • EVA-CLIP (Sun et al., 2023) aligns visual and textual features using contrastive loss, leveraging over 2 billion web image-text pairs and advanced optimization techniques.
  • InternVL3 (Zhu et al., 2025) adopts a unified pre-training approach over both multimodal and pure-text data, enhanced by variable visual position encoding (V2PE) and advanced post-training strategies for improved scalability and effectiveness.
  • InternVL (Chen et al., 2024) leverages contrastive, matching, and generative losses in a multi-stage training process, with a large-scale vision encoder and over 6 billion image-text pairs to align visual and textual representation.
  • BLIP-2 (Li et al., 2023) bridges the modality gap between frozen image encoders and LLMs using a lightweight Q-Former, leveraging contrastive, matching, and generative loss in a two-stage pre-training process over 129 million data with fewer trainable parameters.
  • Qwen2.5-VL (Bai et al., 2025) combines dynamic-resolution Vision Transformer with Window Attention to reduce computational cost while preserving native image resolution.
  • Qwen-VL (Bai et al., 2023) employs a three-stage training process with generative loss, using a VL adapter to align visual and textual features while reducing computational cost over 1.4 billion image-text pairs.
  • CoCa (Yu et al., 2022) adopts task-specific attentional pooling to tailor visual representations for different training objectives, applying contrastive loss to train the first half of the decoder and generative loss to train the full decoder in an end-to-end manner over 5 billion image-text pairs.
  • BEIT3 (Wang et al., 2023) treats images as a foreign language, leveraging a mask-then-predict objective over 36 million image-text pairs to unify vision and language pretraining, and introduces a multiway transformer architecture for general-purpose modeling.
  • LLaVA (Liu et al., 2024a) aligns visual and textual features using a simple MLP with generative loss, leveraging 1.2 million GPT-4 (OpenAI, 2023) generated multimodal instruction-following data for training.
  • DINOv2 (Oquab et al., 2023) uses a self-supervised learning approach, leveraging knowledge distillation and a mask-then-predict strategy over 142 million images to train the vision encoder.
@@ C.1 Results of Hierarchical Granularity Recognition
Figure 2 shows InternVL3’s (Zhu et al., 2025) accuracy in answering true/false and multiple-choice questions within hierarchical granularity recognition task on CUB-200-2011 dataset. In Figure 11, we present additional results for GPT-4o (OpenAI, 2023), Gemini-2.0-flash (Gemini Team, 2024), Qwen2.5-VL (Bai et al., 2025), LLaVA (Liu et al., 2024a) and InternVL (Chen et al., 2024) on CUB-200-2011 (Wah et al., 2011) and iNat2021 (Van Horn et al., 2021) datasets. As shown in the experiments, the accuracy of all models decreases as the granularity becomes finer. When the granularity level reaches the finest level, the models struggle to distinguish between closely related species.
@@
Figure 11: Results of GPT-4o (OpenAI, 2023), Gemini-2.0-flash (Gemini Team, 2024), Qwen2.5-VL (Bai et al., 2025), LLaVA (Liu et al., 2024a) and InternVL (Chen et al., 2024) on true/false and multiple-choice questions across different levels of granularity on CUB-200-2011 (Wah et al., 2011) and iNat2021 (Van Horn et al., 2021) dataset. The x-axis denotes the granularity of the recognition questions.
@@ C.2 Results of Knowledge Bias Estimation
In Figure 3, we observe that LLaVA exhibit highly inconsistent recognition abilities across categories. We also conduct experiments with Qwen2.5-VL, GPT-4o, and Gemini-2.0-flash on fine-grained datasets such as Aircraft (Maji et al., 2013), Flowers102 (Nilsback and Zisserman, 2008) and Stanford Dogs (Khosla et al., 2011). As shown in Figure 12 and Figure 13, all LVLMs display similar trends, indicating inconsistent recognition abilities across fine-grained categories. However, after fine-tuned on datasets with balanced occurrences of fine-grained categories, LVLMs demonstrate remarkable recognition abilities across all fine-grained categories.
@@
Figure 13: Comparison of the original and fine-tuned Qwen2.5-VL (Bai et al., 2025) models on occurrence-balanced fine-grained aircraft categories. True/false question accuracy for each category is ranked, with blue dots representing the original model and yellow dots the fine-tuned model.
@@ C.3 Results of Attribute Recognition
Table 10: Attribute recognition accuracy of LLaVA (Liu et al., 2024a) on the CUB-200-2011 (Wah et al., 2011) dataset (values in parentheses represent the average accuracy for each attribute).
@@
Table 11: Attribute recognition accuracy of BLIP2 (Li et al., 2023) on the CUB-200-2011 (Wah et al., 2011) dataset (values in parentheses represent the average accuracy for each attribute).
@@
Table 12: Attribute recognition accuracy of InternVL (Chen et al., 2024) on the CUB-200-2011 (Wah et al., 2011) dataset (values in parentheses represent the average accuracy for each attribute).
@@
Table 13: Attribute recognition accuracy of Gemini-2.0-flash (Gemini Team, 2024) on the CUB-200-2011 (Wah et al., 2011) dataset (values in parentheses represent the average accuracy for each attribute).
@@
Table 14: Attribute recognition accuracy of Qwen2.5-VL (Bai et al., 2025) on the CUB-200-2011 (Wah et al., 2011) dataset (values in parentheses represent the average accuracy for each attribute).
@@ D.1 Qualitative Analysis of Granularity Inconsistency in LVLM Alignment Data
In practice, ensuring fully consistent fine-grained granularity across all image-text pairs is often infeasible, especially when relying on web-scale or weakly labeled data. In our retraining experiment in Table 4, we made efforts to construct more consistent alignment data, but some residual mismatch may still exist.
@@
D.2 Improving the fine-grained discriminability of visual features during the alignment stage can enhance LVLM performance on fine-grained tasks.
In Table 4, we can find that the alignment strategy might impair the fine-grained discriminability of visual features. We then conduct further analysis and find that improving the fine-grained discriminability of visual features during the alignment stage can enhance LVLM performance on fine-grained tasks.
Specifically, we compare the two variants of LLaVA from Table 4 on fine-grained short-answer questions: (1) Vanilla LLaVA, where the vision-language alignment is trained on image-text pairs with granularity inconsistencies, (2) Retrained LLaVA, where the alignment module is trained on data with matched granularity.
The results in Table 15 show that Retrained LLaVA consistently outperforms Vanilla LLaVA over all datasets, indicating that improving the fine-grained discriminability of visual features during the alignment stage can enhance LVLM performance on fine-grained tasks.
Building on this finding, we believe that incorporating contrastive learning objectives (e.g., patch- or region-level contrastive loss) during the alignment stage may further help preserve discriminative visual information.
Table 15: Linear prob classification results of LLaVA visual features and fine-tuned results of two variants of LLaVA on fine-grained short asnwer questions.
Datasets Linear Fine-tuned
Origin Aligned Aligned-FG Vanilla LLaVA Retrained LLaVA
CUB-200-2011 79.77 73.17 75.06 85.60 86.32
Stanford Dogs 81.24 78.14 80.69 86.49 87.58
Stanford Cars 87.57 83.90 85.63 90.55 91.73
Food-101 94.27 93.35 94.32 95.25 95.74
D.3 Results of Classification Across Multi-categories
@@
Figure 15: Classification results of LVLM visual features on fine-grained datasets. “Single” denotes accuracy from training on a single meta-category, while “Multiple” reflects accuracy from training on a unified dataset combining multiple meta-categories.
Appendix E LLM Usage Disclosure
We acknowledge the use of a Large Language Model (LLM) as a general-purpose assistive tool during the preparation of this paper. Specifically, the LLM was employed in a limited capacity to refine the language, improve clarity, and polish the writing. The model did not play a significant role in research ideation, conceptualization, experimental design, or substantive content generation. All core ideas, analyses, and contributions are entirely the work of the authors.